Miriam was so happy that she is the first honor and she cannot wait to tell her parents about it. So, she ran from SCNHS Campus to her house which is 4.6 km north away. Upon reaching her house, her mother was in their hacienda which is 150 meters north away from their house, so she ran again until she reached her mother. What is the total distance and displacement of Miriam in meters? (Please show the solution for convention)

1
Expert's answer
2021-11-16T09:59:35-0500

The total distance:


"D=4600+150=4750 \text{ m}."

The total displacement is the difference between the initial and final points. Since Miriam run along the straight line forward, the displacement equals the distance:


"d=D=4600+150=4750\text{ m}."